Construct the confidence interval for the ratio of the population variances given the following sample statistics. Round your answers to four decimal places. n1=12 , n2=22, s12=92.101, s22=51.453, 95% level of confidence
Answer #1

Answer:

Given,

n1 = 12

n2 = 22

s12 = 92.101

s22 = 51.453

degree of freedom = n1 - 1

= 12 -1

df1 = 11

df2 = n2 - 1

= 22 - 1

= 20

(s1^2/s2^2) / F(alpha/2 , 11 , 20) < \sigma _{1}^{2} / \sigma _{2}^{2} < (s1^2/s2^2) / F(1 - alpha/2 , 11 , 20)

= 92.101/51.453 / 2.7209 < \sigma _{1}^{2} / \sigma _{2}^{2} < 92.101/51.453 / 0.3100

= 0.6579 < \sigma _{1}^{2} / \sigma _{2}^{2} < 5.7742